OpenStudy (anonymous):

HELP! Write the point-slope form of the equation of the line through the given points. through: (-3, -1) and (2, 4) Question 23 options: A.) y + 1 = -(x + 3) B.) y – 3 = x + 1 C.) y + 1 = x + 3 D.) y + 3 = x + 1

OpenStudy (anonymous):

I thought it was B

OpenStudy (anonymous):

\[y-y _{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\left( x-x _{1} \right)\]

OpenStudy (anonymous):

you are not correct.

OpenStudy (anonymous):

it should start by y-(-1)i.e.,y+1 or y-4

OpenStudy (anonymous):

How would you get the x part?

OpenStudy (anonymous):

either x-(-3) i.e.,x+3 or x-2

OpenStudy (anonymous):

then find the slope as given above.

OpenStudy (anonymous):

So it's C?

OpenStudy (anonymous):

it is a guess or calculated

OpenStudy (anonymous):

anyway you are correct.